OpenStudy (anonymous):

simplify square root of 63

OpenStudy (asnaseer):

use the fact that:\[63 = 7 * 9=7*3^2\]

OpenStudy (anonymous):

i need to give an answer

OpenStudy (anonymous):

do i just put 7.94?

OpenStudy (asnaseer):

so take the square root of the expression I just gave you without actually working out the square root of 7

OpenStudy (anonymous):

7.94 is not a simplified answer.. it's a calculator approximation.

OpenStudy (asnaseer):

HINT:\[\sqrt{x*y^2}=y\sqrt{x}\]

OpenStudy (anonymous):

3 \[3\sqrt{7}\]

OpenStudy (asnaseer):

thats it - perfect!

OpenStudy (asnaseer):

(your 2nd line)

OpenStudy (asnaseer):

please post each question separately

OpenStudy (asnaseer):

you should close this one and post a new one